OpenStudy (anonymous):

Find the equation of the line that passes through the points (0.5, 6.25) and (3, -1.25). Write your answer in Standard Form.

OpenStudy (anonymous):

@Janina Read the previous posts, learn how to find slope, can you?

OpenStudy (ghazi):

(y-6.25))=(-1.25-6.25)/(3-0.5)*(x-0.5) simplify it

OpenStudy (ghazi):

\[(y-6.25)=\frac{ -7.5 }{ 2.5 }*(x-0.5)=-3(x-0.5)\]